# FLARE: Fast Low-rank Attention Routing Engine

### v2

@@ A.1 Input/ output projection
We implement a residual MLP block to serve as a flexible non-linear function approximator. Given input/output dimensions $C_{\text{i}}$ and $C_{\text{o}}$, the layer first applies a linear transformation to a hidden space of size $C_{\text{h}}$, followed by $L$ residual layers, each consisting of a linear layer with GELU activation (Hendrycks & Gimpel, 2016). These are the only instances of pointwise nonlinear activations in the model. An optional input residual connection is applied after the first layer when $C_{\text{i}}=C_{\text{h}}$, and an optional output residual connection is applied at the end when $C_{\text{h}}=C_{\text{o}}$. The final output is projected to dimension $C_{\text{o}}$ via a linear layer. This design allows control over depth and expressivity while preserving stability through residual connections.
@@ A.2 FLARE block
Appendix B Spectral analysis
1import torch.nn.functional as F
2def flare_multihead_mixer_inefficient(q, k, v):
3
4 # Args - q: [H, M, D], k, v: [B, H, N, D]
5 # Ret - y: [B, H, N, D]
6
7 # Compute projection weights
8 scores = q @ k.mT # [B H M N]
9 W_encode = F.softmax(scores, dim=-1) # [B H M N]
10 W_decode = F.softmax(scores.mT, dim=-1) # [B H N M]
11
12 # Encode: Project to latent sequence (M tokens)
13 z = W_encode @ v
14
15 # Decode: Project back to input space (N tokens)
16 y = W_decode @ z
17
18 return y
@@ Appendix B Spectral analysis
B.1 Eigenanalysis procedure
We exploit the low-rank structure of the global communication matrix $W=W_{\text{decode}}\cdot W_{\text{encode}}\in\mathbb{R}^{N\times N}$ to obtain its eigen decomposition in $\mathcal{O}(M^{3}+M^{2}N)$ time, compared to the $\mathcal{O}(N^{3})$ cost for a dense communication matrix. We find the eigen-decomposition of the FLARE attention matrix without actually forming the $N\times N$ matrix. We first note that $W_{\text{encode}}$ and $W_{\text{decode}}$ can be written in terms of the exponentiated score matrix $A=\exp(Q\cdot K^{T})\in\mathbb{R}^{M\times N}$ as
@@ B.1 Eigenanalysis procedure
Figure 7: $M=64$ nonzero eigenvalues of $W_{h}$ (Eq. 9) for $H=8$ heads in different blocks of FLARE with $B=8$ blocks $C=64$ features trained on the elasticity dataset ($972$ points per input). The number of latents ($M$) is depicted with a dashed red line, and FP32 precision with a dashed black line.
B.2 Qualitative Analysis
@@
Figure 7 presents the $M=64$ nonzero eigenvalue spectra of an FLARE model trained on the elasticity dataset with $M=64$ latents. Some observations are as follows. In all blocks, especially block 1, the eigenvalues drop sharply within the first 102010-20 indices. This indicates that even though the communication matrices $W_{h}$ could have rank up to $M=64$, most of the energy (information) is captured by a much smaller subset of modes. This result validates the hypothesis that the global communication pattern is inherently low-rank.
@@
Appendix C Benchmarking and Comparison
C.1 Benchmark metrics
@@
Benchmark #Dim Grid Type #Points
#Input/ Output
Features
#Train/ Test
Cases
@@
LPBF 1,000–50,000 1100 / 290
C.2 Benchmark datasets
We evaluate all models on six benchmark datasets and our proposed AM dataset, each designed to assess generalization, scalability, and robustness to domain irregularity in PDE surrogate modeling. A summary is provided in Table 3.
@@ C.2 Benchmark datasets
We introduce the laser powder bed fusion (LPBF) additive manufacturing dataset which involves field prediction on complex geometries. In metal additive manufacturing, variations in design geometry can affect the dimensional accuracy of the part and lead to shape distortions. We perform several LPBF process simulations of a set of geometries to obtain the deformation field over the geometry. We select a subsample of the dataset with up to 50,000 points per geometry and divide it into 1,100 training and 290 cases. We train models to learn the $Z$ (vertical) component of the deformation field at each grid point. Additional details are provided in Appendix G.
C.3 Benchmark models and training details
@@
We evaluate Transolver, GNOT with the hyperparameter configurations provided in Wu et al. (2024) and LNO on the ones provided in Wang & Wang (2024a) on the 2D test cases. For the remaining problems, we choose the best performing parameter set from the 2D cases.
Unless otherwise stated, all models are trained with the AdamW (Loshchilov & Hutter, 2019) optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) to minimize the relative $L_{2}$ error. We use the OneCycleLR (Smith & Topin, 2019) scheduler with 10% of epochs dedicated to warming-up to a learning rate of $10^{-3}$, followed by cosine decay. We train on LPBF for 250 epochs, and 500 epochs for all other models. Note that we use gradient clipping with max_norm=1.0=1.0 unless otherwise stated. Unless otherwise stated, the weight decay regularization parameter is set to $10^{-5}$ for the 2D test cases, $10^{-4}$ for DrivAerML-40k, and $10^{-4}$ for LPBF. The batch size is set to 2 for the 2D problems and 1 for the 3D problems unless otherwise stated. Unless otherwise stated, all models are trained in full precision (FP32).
@@
For PerceiverIO, we use $B=8$, $C=128$, $H=8$, and set the latent sequence length to $M=1,024$.
@@
Transolver (Wu et al. (2024)) introduces Physics-Attention: each mesh point is softly assigned to a few learnable slices, shrinking thousands of points to only tens of tokens. Self-attention runs on this compact set and the tokens are then desliced back to points. Following the hyperparameter recommendations in the code of Wu et al. (2024), Transolver is trained with 30% of the steps dedicated to warm-up, and gradient clipping with max_norm =0.1=0.1. The recommended batch size for Transolver is 1 for elasticity, and 4 for the remaining 2D problems. For the 3D problems, we set the batch size to 1.
@@
LNO (Wang & Wang (2024b)) moves computation into a small latent space. An embedder lifts the input field, cross-attention compresses $N$ points into $M$ latent tokens, transformers act solely on these tokens, and a decoder maps them to any query location. In line with the recommended hyperparameter configuration in Wang & Wang (2024b), we set $\beta_{2}=0.99$ in the AdamW optimizer, do warmup for the first 20% of epochs, and clip gradient norms greater than 1,000. The number of hidden features are set to $C=192$ for elasticity, and $C=128$ for all other test cases. The number of residual layers is set to 3 for elasticity and 4 for other test cases. The number of latent self attention blocks is 8 for pipe and airfoil test cases and 4 for all other cases. The number of latent modes is set to 256 for all test cases. The batch size during training is set to 4 for the 2D test cases and 1 for the 3D test cases. The LNO code also recommends weight decay regularization of $5\cdot 10^{-5}$ for the 2D test cases.
In running our experiments, we noticed discrepancies between our LNO results and those presented in their article (Wang & Wang, 2024a). Upon further investigation, we found that the datasets used in the LNO paper and in the original Transolver paper are not the same. For example, in the Elasticity dataset, LNO was trained and tested on a $1000/1000$ split, whereas Transolver used a $1000/200$ split. In the Darcy dataset, LNO employed a higher resolution of $241\times 241$, compared to $85\times 85$ in Transolver. Because these differences make direct comparison unreliable, all models (including LNO and Transolver) were re-trained and evaluated on the standardized Transolver splits and resolutions to ensure fairness.
@@
GNOT (Hao et al. (2023)) employs heterogeneous normalized attention, separately normalizing keys and values, to fuse multiple input fields on irregular meshes. A learnable geometric gate decomposes the domain and routes tokens to scale-specific expert MLPs. Linear cost attention plus this gating scales to large problems and surpasses earlier operator learners.
Following the hyperparameter recommendations outlined in the code of Wu et al. (2024), GNOT is trained with 30% of the steps dedicated to warm-up, and gradient clipping with max_norm = $0.1$. The recommended batch size for Transolver is 1 for elasticity, and 4 for the remaining 2D problems. The batch size is set to 2 for elasticity, 4 for the remaining 2D benchmarks and 1 for the 3D benchmarks.
@@
For all problems, we employ $B=8$ blocks with a feature dimension of $C=64$. We set the number of residual layers and the number of key/value projection layers to 3, and vary the head dimension as $D\in\{4,8\}$, the number of latent tokens as $M\in\{64,128,256\}$ The hyperparameters for each test case are presented in Table 5.
@@
Appendix D Field-prediction on million-point geometries
All FLARE models in this study are trained in mixed precision to take advantage of the flash attention algorithm. We train for 500 epochs with the OneCycleLR scheduler (Smith & Topin, 2019) where the first $5\%$ of epochs are spent warming up to a learning rate of $5\cdot 10^{-4}$ followed by cosine decay.
Appendix E Discussion on the design principles of FLARE.
@@ Appendix D Field-prediction on million-point geometries
the output token $y_{n}$ only receives substantial information from the latents whose query pattern matches its key $k_{n}$ strongly. In this sense, each latent acts as both a selective pooling hub and broadcaster, routing information back along the pathways that originally recruited it. Together, these gather–scatter stages form a low-rank, butterfly-like transformation: first contracting $N$ tokens into $M\ll N$ global features, and then expanding back to $N$ outputs, enabling efficient, yet expressive, global communication.
@@
Appendix F Model analysis and ablations
@@
Number of heads.
A key hypothesis behind our design is multiple simultaneous (head-wise parallel) low-rank projections collectively approximate a full attention map. Thus, for a fixed total feature dimension ($C$), using a smaller head dimension ($D$) should improve performance by allowing more such parallel projections. To test this, we evaluate FLARE on the elasticity dataset with varying head counts while keeping the total feature dimension $C=64$ and number of blocks $B=8$ fixed. As shown in Figure 10, FLARE achieves the lowest test error when $D=4$ or $D=8$, in contrast to standard transformers which often use D=16D=163232. Notably, because our head dimensions are smaller, we use a scaling factor of $1$ instead of the typical $\sqrt{D}$ employed in transformers (Vaswani et al., 2017), as the dot product magnitudes are sufficiently small without additional scaling.
Appendix G Benchmark dataset of additive manufacturing simulations

### v3

@@ B.1 Input/ output projection
We implement a residual MLP block to serve as a flexible non-linear function approximator. Given input/output dimensions $C_{\text{i}}$ and $C_{\text{o}}$, the layer first applies a linear transformation to a hidden space of size $C_{\text{h}}$, followed by $L$ residual layers, each consisting of a linear layer with GELU activation (Hendrycks and Gimpel, 2016). These are the only instances of pointwise nonlinear activations in the model. An optional input residual connection is applied after the first layer when $C_{\text{i}}=C_{\text{h}}$, and an optional output residual connection is applied at the end when $C_{\text{h}}=C_{\text{o}}$. The final output is projected to dimension $C_{\text{o}}$ via a linear layer. This design allows control over depth and expressivity while preserving stability through residual connections.
@@ B.2 FLARE block
Appendix C Spectral analysis
import torch.nn.functional as F
def flare_multihead_mixer_inefficient(Q, K, V):
# Args - Q: [H M D], K, V: [B H N D]
# Ret - Y: [B H N D]
# Compute projection weights
scores = Q @ K.mT # [B H M N]
W_encode = F.softmax(scores, dim=-1) # [B H M N]
W_decode = F.softmax(scores.mT, dim=-1) # [B H N M]
# Encode: Project to latent sequence (M tokens)
Z = W_encode @ V
# Decode: Project back to input space (N tokens)
Y = W_decode @ Z
return Y
@@ Appendix C Spectral analysis
C.1 Eigenanalysis procedure
We exploit the low-rank structure of the global communication matrix $W=W_{\text{decode}}\cdot W_{\text{encode}}\in\mathbb{R}^{N\times N}$ to obtain its eigendecomposition in $\mathcal{O}(M^{3}+M^{2}N)$ time, compared to the $\mathcal{O}(N^{3})$ cost for a dense communication matrix. We find the eigendecomposition of the FLARE attention matrix without actually forming the $N\times N$ matrix. We first note that $W_{\text{encode}}$ and $W_{\text{decode}}$ can be written in terms of the exponentiated score matrix $A=\exp(Q\cdot K^{T})\in\mathbb{R}^{M\times N}$ as
@@ C.1 Eigenanalysis procedure
C.2 Qualitative Analysis
@@
Figure 12 (middle) presents the $M=64$ nonzero eigenvalue spectra of an FLARE model trained on the elasticity dataset with $M=64$ latents. Some observations are as follows. In all blocks, especially block 1, the eigenvalues drop sharply within the first 102010-20 indices. This indicates that even though the communication matrices $W_{h}$ could have rank up to $M=64$, most of the energy (information) is captured by a much smaller subset of modes. This result validates the hypothesis that the global communication pattern is inherently low-rank.
@@
Appendix D Benchmarking and Comparison
D.1 Benchmark metrics
@@
Benchmark #Dim Grid Type #Points #Input/ Output Features #Train/ Test Cases
@@
LPBF 3D Unstructured 1,000–50,000 3 / 1 1100 / 290
D.2 Benchmark datasets
We evaluate all models on five benchmark datasets and our proposed AM dataset, each designed to assess generalization, scalability, and robustness to domain irregularity in PDE surrogate modeling. A summary is provided in Table 3.
@@ D.2 Benchmark datasets
We introduce the laser powder bed fusion (LPBF) additive manufacturing dataset which involves field prediction on complex geometries. In metal additive manufacturing, variations in design geometry can affect the dimensional accuracy of the part and lead to shape distortions. We perform several LPBF process simulations of a set of geometries to obtain the deformation field over the geometry. We select a subsample of the dataset with up to 50,000 points per geometry and divide it into 1,100 training cases and 290 test cases. We train models to learn the $Z$ (vertical) component of the deformation field at each grid point. Additional details are provided in Appendix H.
D.3 Benchmark models and training details
@@
We evaluate Transolver, GNOT with the hyperparameter configurations provided in (Wu et al., 2024) and LNO on the ones provided in (Wang and Wang, 2024b) on the 2D test cases. For the remaining problems, we choose the best performing parameter set from the 2D cases.
Unless otherwise stated, all models are trained with the AdamW (Loshchilov and Hutter, 2019) optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) to minimize the relative $L_{2}$ error. We use the OneCycleLR (Smith and Topin, 2019) scheduler with 10% of epochs dedicated to warming-up to a learning rate of $10^{-3}$, followed by cosine decay. We train on LPBF for 250 epochs, and 500 epochs for all other models. Note that we use gradient clipping with max_norm=1.0=1.0 unless otherwise stated. Unless otherwise stated, the weight decay regularization parameter is set to $10^{-5}$ for the 2D test cases, $10^{-4}$ for DrivAerML-40k, and $10^{-4}$ for LPBF. The batch size is set to 2 for the 2D problems and 1 for the 3D problems unless otherwise stated. Unless otherwise stated, all models are trained in full precision (FP32).
@@
For PerceiverIO, we use $B=8$, $C=128$, $H=8$, and set the latent sequence length to $M=512$ for all test cases.
@@
Transolver ((Wu et al., 2024)) introduces Physics-Attention: each mesh point is softly assigned to a few learnable slices, shrinking thousands of points to only tens of tokens. Self-attention runs on this compact set and the tokens are then desliced back to points. Following the hyperparameter recommendations in the code of (Wu et al., 2024), Transolver is trained with 30% of the steps dedicated to warm-up, and gradient clipping with max_norm =0.1=0.1. The recommended batch size for Transolver is 1 for elasticity, and 4 for the remaining 2D problems. For the 3D problems, we set the batch size to 1.
Transolver++.
We evaluated Transolver++ using the latest version of the official code release (https://github.com/thuml/Transolver_plus) together with the hyperparameters reported in the paper. All experiments were run using PyTorch 2.7 and CUDA 12.8; the supplementary material provides a script to instantiate the standard environment used for all benchmark runs.
@@
LNO ((Wang and Wang, 2024a)) moves computation into a small latent space. An embedder lifts the input field, cross-attention compresses $N$ points into $M$ latent tokens, transformers act solely on these tokens, and a decoder maps them to any query location. In line with the recommended hyperparameter configuration in (Wang and Wang, 2024a), we set $\beta_{2}=0.99$ in the AdamW optimizer, do warmup for the first 20% of epochs, and clip gradient norms greater than 1,000. The number of hidden features are set to $C=192$ for elasticity, and $C=128$ for all other test cases. The number of residual layers is set to 3 for elasticity and 4 for other test cases. The number of latent self attention blocks is 8 for pipe and airfoil test cases and 4 for all other cases. The number of latent modes is set to 256 for all test cases. The batch size during training is set to 4 for the 2D test cases and 1 for the 3D test cases. The LNO code also recommends weight decay regularization of $5\cdot 10^{-5}$ for the 2D test cases.
Through communication with the LNO authors (see https://github.com/L-I-M-I-T/LatentNeuralOperator/pull/6), we learned that the datasets used in the LNO paper and in the original Transolver paper are not the same.
In running our experiments, we noticed discrepancies between our LNO results and those presented in their article (Wang and Wang, 2024b). Upon further investigation, we found that the datasets used in the LNO paper and in the original Transolver paper are not the same. For example, in the Elasticity dataset, LNO was trained and tested on a $1000/1000$ split, whereas Transolver used a $1000/200$ split. In the Darcy dataset, LNO employed a higher resolution of $241\times 241$, compared to $85\times 85$ in Transolver. Because these differences make direct comparison unreliable, all models (including LNO and Transolver) were re-trained and evaluated on the standardized Transolver splits and resolutions to ensure fairness.
@@
GNOT ((Hao et al., 2023)) employs heterogeneous normalized attention, separately normalizing keys and values, to fuse multiple input fields on irregular meshes. A learnable geometric gate decomposes the domain and routes tokens to scale-specific expert MLPs. Linear cost attention plus this gating scales to large problems and surpasses earlier operator learners.
Following the hyperparameter recommendations outlined in the code of (Wu et al., 2024), GNOT is trained with 30% of the steps dedicated to warm-up, and gradient clipping with max_norm = $0.1$. The recommended batch size for Transolver is 1 for elasticity, and 4 for the remaining 2D problems. The batch size is set to 2 for elasticity, 4 for the remaining 2D benchmarks and 1 for the 3D benchmarks.
@@
All attention operations in FLARE are implemented using scaled_dot_product_attention (SDPA) in torch.nn.functional, as shown in Figure 3; this allows PyTorch to dispatch to fused SDPA backends (e.g., FlashAttention-style kernels) when available. For all problems, we employ $B=8$ blocks with a feature dimension of $C=64$. We set the number of residual layers and the number of key/value projection layers to 3, and vary the head dimension as $D\in\{4,8\}$, the number of latent tokens as $M\in\{64,128,256\}$ The hyperparameters for each test case are presented in Table 5.
@@
Appendix E Field-prediction on million-point geometries
Experimental setup.
We train FLARE on the DrivAerML dataset (Ashton et al., 2024) using meshes subsampled to contain $N=10^{6}$ points per geometry. All experiments are run on a single Nvidia H100 80GB GPU with batch size $1$. Unless otherwise stated, models use feature dimension $C=64$ and $H=8$ attention heads. We vary the number of latent tokens $M\in\{128,512,1024,2048\}$ and the number of FLARE blocks $B$ (as reported in Figure 5).
Training details.
All FLARE models in this study are trained in mixed precision to take advantage of fused SDPA backends (e.g., FlashAttention-style kernels). We train for 500 epochs with the OneCycleLR scheduler (Smith and Topin, 2019) where the first $5\%$ of epochs are spent warming up to a learning rate of $5\cdot 10^{-4}$ followed by cosine decay. No hidden sparsity, memory offloading, or distributed training is used.
Appendix F Discussion on the design principles of FLARE.
@@ Appendix E Field-prediction on million-point geometries
the output token $y_{n}$ only receives substantial information from the latents whose query pattern matches its key $k_{n}$ strongly. In this sense, each latent acts as both a selective pooling hub and broadcaster, routing information back along the pathways that originally recruited it. Together, these gather-scatter stages form a low-rank, butterfly-like transformation: first contracting $N$ tokens into $M\ll N$ global features, and then expanding back to $N$ outputs, enabling efficient, yet expressive, global communication.
@@
Appendix G Model analysis and ablations
Figure 9: Effect of number of blocks ($B$) and number of latent tokens ($M$) on test accuracy on the Elasticity (left) and Darcy (right) test cases. Experiment details are presented in Appendix G.
Time and memory complexity.
Although vanilla self-attention has the lowest memory cost thanks to the FlashAttention algorithm, which eliminates the need to materialize the score matrices ($Q_{h}\cdot K_{h}^{T}$), its compute time still scales poorly with the sequence length. In contrast, the compute time for FLARE exhibits strong scaling with sequence length. Its memory requirement is marginally greater than vanilla attention due to the presence of deep residual networks for key/value projections, and due to the need to materialize $Z_{h}$, the latent sequence of $M$ tokens. As these costs are marginal compared to the SDPA operation, the curves for different $M$ values of FLARE are somewhat overlapping. Finally, the compute time for Physics Attention of Transolver (Wu et al., 2024) exhibits somewhat good scaling. However, its memory cost and compute time blow up for large slice counts due to the need for materializing the projection matrices.
Number of blocks ($B$) and latent tokens ($M$).
@@ Appendix G Model analysis and ablations
Figure 11: Ablation on the number of latent-space self-attention blocks ($L_{B}$) versus the number of FLARE encode-decode blocks ($B$). Each cell reports: (top) test relative error, (middle) parameter count, and (bottom) training time per epoch. The total cost of a model with $B$ FLARE blocks and $L_{B}$ latent blocks is $\mathcal{O}(B\,NM+BL_{B}\,M^{2})$. The results show that increasing the number of latent-space blocks, as done in LNO-style and Perceiver-style models, yields worse accuracy and poorer speed/parameter tradeoffs than allocating compute to additional FLARE blocks. The optimal regime lies near the top-right corner: many encode-decode blocks and zero latent-space blocks.
Blocks Shared latents Independent latents
1 20.8 21.9
2 10.8 9.99
4 5.91 5.21
8 4.37 3.85
Figure 12: Ablation of shared vs. independent latent tokens across attention heads. Each plot shows the M=64M{=}64 nonzero eigenvalues of the head-specific communication matrices $W_{h}$ (Eq. 9) for FLARE with B=8B{=}8 blocks and C=64C{=}64 features trained on the elasticity dataset. When heads share a single latent sequence (top row), the eigenvalue spectra across heads are nearly identical, indicating similar learned low-rank subspaces. When heads use independent latent slices (bottom row), the eigenvalue decay varies noticeably across heads, reflecting more diverse low-rank structures. The accompanying table reports test relative errors for different depths $B$, showing that models with independent latents consistently achieve lower error.
@@
Effect of head dimension and parallel low-rank projections.
[…]
To examine this hypothesis quantitatively, we vary the number of heads $H$ while keeping the total feature dimension $C=64$ and number of blocks $B=8$ fixed, thereby trading off head dimension $D=C/H$ against the number of parallel low-rank projections available to the model. As shown in Figure 13, FLARE consistently achieves the best accuracy for $D=4$ or $D=8$, outperforming configurations with larger head dimensions. This behavior is the reverse of what is commonly observed in standard transformers—where typical head dimensions are D=16D=163232, but is entirely consistent with FLARE’s architectural role for each head. Larger $D$ increases the per-head representational capacity but reduces the number of parallel low-rank factors; smaller $D$ yields more distinct heads, and thereby more distinct projection–reconstruction pathways, which better approximate a full attention map.
We additionally note a practical implication of this analysis: because the dot-product magnitudes are naturally small for $D\in\{4,8\}$, we use a scaling factor of $1$ rather than the usual $1/\sqrt{D}$ in these cases, simplifying implementation without sacrificing stability. Overall, the head-dimension ablation strongly supports our main architectural claim: FLARE benefits from multiple parallel low-rank projections, and enabling per-head independence is crucial for capturing diverse communication patterns.
Ablation on latent-space blocks vs. FLARE blocks.
To directly investigate whether FLARE’s encode-decode mechanism is responsible for the observed performance gains, we conduct a controlled ablation varying the number of latent-space self-attention blocks ($L_{B}$) and the number of full FLARE blocks ($B$). This experiment spans the continuum between a Perceiver/LNO-like architecture (few encode-decode operations and many latent-space blocks; bottom-left of Figure 11) and a FLARE-like architecture (many encode-decode blocks and no latent-space self-attention; top-right).
Across all parameter budgets, we observe a consistent trend: increasing the number of latent-space blocks degrades accuracy while increasing computational cost. Latent-space self-attention contributes additional parameters and a $\mathcal{O}(M^{2})$ cost per block, yet does not improve the model’s ability to capture global structure. In contrast, allocating the same compute to additional encode-decode blocks, which perform low-rank global mixing via cross-attention, steadily improves accuracy. The lowest errors in the entire grid occur in the top-right corner, corresponding to zero latent-space blocks and the largest number of FLARE blocks.
This ablation provides direct causal evidence supporting FLARE’s architectural choice: global communication is most effectively learned by repeating the low-rank projection/unprojection pathway, not by performing nonlinear refinements inside the latent space. If one interprets the Perceiver $\leftrightarrow$ FLARE spectrum as trading latent refinement for repeated global mixing, our results show that the optimal regime is decisively FLARE-like. For a fixed parameter budget and wall-clock time, the best strategy is to reduce or eliminate latent-space attention and increase the number of encode-decode blocks, validating the design principle that token mixing should occur through repeated low-rank attention projections rather than deeper latent-space transformers.
Ablation on shared vs. independent per-head latent tokens.
To assess the effect of head-wise latent diversity—central to FLARE’s design, we compare two variants: (i) shared-latent models, where all heads use the same latent sequence, and (ii) independent-latent models, where each head receives its own slice of the latent tokens. This isolates whether FLARE’s expressivity arises merely from the low-rank encode-decode structure or whether the ability of different heads to learn distinct low-rank projections provides additional modeling capacity.
The spectral plots in Figure 12 reveal a clear difference between the two settings. With shared latents, all heads exhibit nearly identical eigenvalue decay, implying that they compress and propagate information in similar ways. In contrast, independent latents produce noticeably different spectra across heads, particularly in deeper blocks, indicating that different heads discover complementary low-rank subspaces of the token-to-token communication structure. This diversity is modest in early layers but becomes increasingly pronounced as depth increases.
Crucially, the quantitative results mirror this qualitative behavior: across all depths, models with independent per-head latents achieve lower test relative error despite having comparable parameter counts. This provides causal evidence—not just descriptive visualization—that head-wise independence is beneficial. It supports our claim that FLARE’s mixture of head-specific low-rank projections yields richer communication pathways than architectures that use a single shared projection (e.g., Transolver) or a single latent transformer (e.g., LNO, PerceiverIO). Finally, we note that applying the same spectral analysis to Transolver or LNO is not directly meaningful, since their latent-space transformations are nonlinear and not expressible as a single linear low-rank operator; FLARE’s structure uniquely enables such eigenanalysis.
Appendix H Benchmark dataset of additive manufacturing simulations